OpenStudy (anonymous):

If a line crosses the y-axis at (0, -4) and has a slope of -2, what is the equation of the line?

OpenStudy (ivettef365):

you already have the information you need m = -2 and b = -4 just substitute on the equation of the line which is y = mx +b

OpenStudy (anonymous):

use the point-slope formula \[(y-y_1) = m(x-x_1)\] where m is the slope and (x_1, y_1) is the point